OpenStudy (anonymous):

Use the following formula: lw+- (w+l)e, where e is the absolute error in each length measurement, w is the width, l is the length to calculate the area of a 13.8m x 7.3m plot. What is the absolute error in this case? Somebody show me the working out please??

OpenStudy (anonymous):

A little confusing, but I think this is what I'm looking at: \(\large Area = L*W\pm (W+L)*e\) \(\large Length, L=13.8m\) \(\large Width, W=7.3m\) And you want to solve for e?

OpenStudy (anonymous):

No, I have to find the area, and e is the absolute error but I don't know what the absolute error is. According to my teacher: A measurement is accurate to \[\pm \frac{ 1 }{ 2 }\] of the smallest division on the scale. But I'm not sure how to work with that.

OpenStudy (anonymous):

Ah, ok. See that the measurements are precise to 0.1m, that means the absolute error is ±0.05m

OpenStudy (anonymous):

I'm so stupid, now it all makes sense! Thank you so much!

OpenStudy (anonymous):

No prob' You're welcome.